OpenStudy (anonymous):

easy medal 6+ v/8=4/7

OpenStudy (anonymous):

Solving for v?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

subtract both sides by 6 and then multiply both sides by 8

OpenStudy (anonymous):

this will let you isolate v

OpenStudy (anonymous):

\[6+\frac{ v }{ 8 }=\frac{ 4 }{ 7 }\] \[\frac{ v }{ 8 }=\frac{ 4 }{ 7 }-6\] \[\frac{ v }{ 8 }=\frac{ -38 }{ 7 }\] \[v=\frac{ -304 }{ 7 }\]

OpenStudy (anonymous):

or we can write this as an improper fraction. \[v=-43\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

: ) np